OpenStudy (anonymous):

Fill in the blank to make the expression a perfect square. v^2 + 8v + ___

OpenStudy (anonymous):

two steps what is half o f 8?

OpenStudy (anonymous):

...

OpenStudy (welshfella):

you want (v + a)v + a) = v^2 + 8v + a^2 and 2av must equal 8v so what is a equal to?

OpenStudy (idealist10):

(8/2)^2=4^2=16 For perfect square problems, you just divide by 2 for your second term and just square it.

OpenStudy (anonymous):

Thanks! :-)

OpenStudy (idealist10):

You're welcome!